OpenStudy (vanillaangelic):

Please explain this. Find the volume of a cylinder with a diameter 5 inches and height 5 feet. (Pi = 22/7)

OpenStudy (shamim):

Make clearer ur question plz!

OpenStudy (mathstudent55):

Use this formula: \(\Large V_{cylinder} = \dfrac{1}{3} \pi r^2 h\) where r = radius of the base, and h = height of the cylinder.

OpenStudy (mathstudent55):

Also, I think you means to use pi = 22/7, not that the radius is 22/7.

OpenStudy (shamim):

I think volume of a cylinder=pi*r^2*h @mathstudent55

OpenStudy (mathstudent55):

LOL. I've been online too long. Yes, of course you're right. My formula above is for a cone.

OpenStudy (mathstudent55):

\(\Large V_{cylinder} = \pi r^2h\)

OpenStudy (vanillaangelic):

Mind if you explain how do it step by step? Thanks. I really don't understand.

OpenStudy (mathstudent55):

Ok. Here is the explanation. For the formula you need the radius of the base of the cylinder. You were given the diameter. Do you know the relationship between the diameter and the radius of a circle?

OpenStudy (vanillaangelic):

No..

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

A chord of a circle is a line segment that has as its endpoints any two points on the circle. A chord that passes through the center of a circle is a diameter. A radius of a circle is a line segment that has the center as an endpoint and a point on the circle as its other endpoint.

OpenStudy (vanillaangelic):

Ok.

OpenStudy (mathstudent55):